OpenStudy (anonymous):

3 5/9 = improper fraction?

OpenStudy (lgbasallote):

improper fraction looks like \[\large \frac{\text{big number}}{\text{small number}}\]

OpenStudy (lgbasallote):

i.e. \[\large \frac{22}{7} , \frac 73, \frac 52, \frac 83, \text{etc.}\]

OpenStudy (anonymous):

yes, but how do i change it?

OpenStudy (lgbasallote):

oh..i thought you were asking if it's improper

OpenStudy (lgbasallote):

what's your denominator?

OpenStudy (anonymous):

9

OpenStudy (lgbasallote):

and what's your whole number?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3 5/9

OpenStudy (lgbasallote):

nope that's the numerator

OpenStudy (lgbasallote):

3 5/9 is the mixed fraction

OpenStudy (lgbasallote):

the whole number is the number outside of the fraction

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

oops wait

OpenStudy (lgbasallote):

it's 3 sorry

OpenStudy (lgbasallote):

so multiply that with 9

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

what's the product?

OpenStudy (anonymous):

27

OpenStudy (lgbasallote):

now ADD it to the numerator

OpenStudy (anonymous):

32

OpenStudy (lgbasallote):

so your improper fraction is \[\frac{32}{5}\]

OpenStudy (lgbasallote):

\[\frac{32}{9}\] sorry

OpenStudy (anonymous):

ok